OpenStudy (anonymous):

how do you graph the function rule y=2x+4

OpenStudy (anonymous):

line with slope 2 and y intercept 4. means crosses y - axis at (0,4) and every time you go to the right one you go up 2

OpenStudy (anonymous):

plot some points and connect the dots with a ruler.

OpenStudy (anonymous):

whats a slope

OpenStudy (anonymous):

if you graph the line you will see that every every time x increases by 1, y increases by 2. for example on you line you have the point (1,6) because if x = 1 then 7 = 2*1+4=6 then if you increase x by 1 unit to get 2, y will be 2*2+4= 8, so you will have the point (2,8) on the graph

OpenStudy (anonymous):

x goes over 1 unit, y goes up 2 units.

OpenStudy (anonymous):

i meant y = 2*1+4 not 7 = 2*1+4 sorry